Tohban Report 2013-11-06

From RHESSI Wiki

Jump to: navigation, search

Tohban Reports
Start Date: 30 Oct 2013
End Date: 6 Nov 2013
Tohban: Hazel Bain
Tohban email:
Next Tohban: Nicole Duncan
List all reports


Solar Activity

Solar activity this week was high with 6 M class and 1 X class flare! The main flaring activity was produced by ARs 11884, at the start of the week and now by AR 11890, which is currently classified as a beta/gamma/delta region and produced an X3 class flare yesterday. AR 11890 is still in the eastern hemisphere and is expected to continue producing flares. In addition, there are 6 other ARs on the disk, several of which are producing lower C class flares. STEREO B shows a few more ARs will rotate onto the disk within the next few days.

RHESSI observed the solar eclipse during the 3 orbits. The eclipse lasted ~10 mins during each orbit.

How many GOES flares occurred?

 Flares above B, C, M, X class were      3    46     6     1

And how many of these are listed in the RHESSI flare list?

 Flares above B, C, M, X class were      1    36     3     0

And how many had EXCELLENT coverage?

 Flares above B, C, M, X class were      1    17     2     0

There were RHESSI flares/GOES flares 404 / 56 over the time range 30-Oct-13 06-Nov-13

Memory Management

Due to high solar activity, the SSR fill reached a maximum of 31% but is now emptying at the end of the daily pass set. We remained in Active / Vigorous mode this week and with front and rear eclipse data turned off (Plus/minus 4 minute data is still on for both fronts and rear.)

Spacecraft Status

Cold plate 2 is currently at 114 K.

Data Gaps

There were a few 5 min data gaps due to malformed data packets on the 31st and 1st.

Detector issues

Decimation Active/Vigorous
HLAT Decimation Rear decimation weight 6, no front decimation
Night time data (fronts) Disabled, except for +/- 4 min
Night time data (rears) Disabled, except for +/- 4 min (Changed enabled->disabled this week)
Require extra passes? No
Requirement for moving pointer? No
Attenuator operation Normal
Detector problems? None
Personal tools